# Hierarchies of direct powers, ultrapowers and cumulative powers

### 16 Mar 2026 v1

@@ 2.2. Preservation and properties of cumulative powers
Therefore, for $n<\omega$, $\mathcal{F}^{+}_{n}(\mathcal{A})$ inherits any first-order property from $\mathcal{A}$ characterized by non-collapsible direct power sentences; and for $\beta\in\mathbf{On}$, $\mathcal{F}_{\beta}(\mathcal{A})$ inherits any first-order property from $\mathcal{A}$ characterized by constant-free non-collapsible direct power sentences – since, for $\beta>0$, $\mathcal{F}_{\beta}(\mathcal{A})$ does not have an interpretation of the constants of $\sigma$. Not only that, those are the precisely the classes of first-order sentences inherited by cumulative powers. In other words, just as the fragment preserved by reduced products is precisely the one given by Horn sentences, and the fragment preserved by direct products is precisely the one characterized by Weinstein, the fragment preserved by (finite) cumulative powers is precisely the one given by non-collapsible direct power sentences.
Notice the restrictions of the results on non-collapsible formulas concern precisely equality. That is related to the definition of the extended functions. In our former example, for $j\in\mathbb{N}$, $\vartheta^{1}_{j}(\overline{3})=\overline{3}(j)=3$, $\vartheta^{1}_{j}(\overline{-1})=-1$ and $\vartheta^{1}_{j}(2)=2$, so $\forall j\in\mathbb{N}\big(\vartheta^{1}_{j}(\overline{3})+\vartheta^{1}_{j}(\overline{-1})=\vartheta^{1}_{j}(2)\big)$. However, $\mathcal{F}_{1}(\mathbb{N})\not\models\overline{3}+\overline{-1}=2$, as $\overline{3}+\overline{-1}=\overline{2}\neq 2$. The restriction has also a deeper meaning. It concerns the difference between the definition of other extended relations and the equality relation. While at each level $\beta$ every other relation is defined by using the $\vartheta$ function, and therefore take into account all of the levels lesser than $\beta$ – so that, for example, if $\rho(a)<\rho(b)=\beta$ for a successor $\beta$, then $\mathrm{R}_{\beta}ab$ iff $\mathrm{R}_{\beta-1}ab(j)$ –, the equality relation is interpreted as real identity, and thus holds between two objects iff they are the same object, regardless of how they are hereditarily related. In a few words, whereas every other relation is hereditarily defined, equality is not.
@@
Let $\mathcal{A}$ and $\mathcal{B}$ be $\sigma$-structures, and their cumulative power hierarchies be respectively generated by the families of index sets $\{I_{\kappa}\}_{\kappa\in\mathbf{On}}$ and $\{J_{\kappa}\}_{\kappa\in\mathbf{On}}$. If $\mathcal{A}\hookrightarrow\mathcal{B}$ and for each $\beta\in\mathbf{On}$ $|I_{\beta}|\leq|J_{\beta}|$, then for any $\beta,\gamma\in\mathbf{On}$ such that $\beta\leq\gamma$, $\mathcal{F}_{\beta}(\mathcal{A})\hookrightarrow\mathcal{F}_{\gamma}(\mathcal{B})$.
@@ 3. Direct powers and the hierarchy
As the past results show, the $\mathcal{F}$-hierarchies generated, as a whole, present a structure that resembles those of their generating structure, but nevertheless fail to instantiate the generating structure itself. How can we recover more of the first-order properties of the generating structures? A way of doing that is by restricting the elements of the hierarchy to those of the parallel hierarchy which is not cumulative – in other words, a hierarchy if direct powers. Consider the following definition.
@@
Let $\{I_{\alpha}\}_{\alpha\in\mathbf{On}}$ be a family of index sets. The direct power hierarchy generated by $\mathcal{A}$ and $\{I_{\alpha}\}_{\alpha\in\mathbf{On}}$ is composed by the structures $\Pi_{\beta}(\mathcal{A})$, where each $\Pi_{\beta^{+}}(\mathcal{A})$ is the direct power of $\Pi_{\beta}(\mathcal{A})$ with index set $I_{\beta}$, and for a limit $\lambda$, $\Pi_{\lambda}(\mathcal{A})$ is the limit of embeddings of the preceding structures. More precisely:
@@
By Corollary 2.15, we know what sentences are preserved by $\Pi_{\beta}(\mathcal{A})$. Suppose we let the same family $\{I_{\alpha}\}_{\alpha\in\mathbf{On}}$ define both a cumulative power and a direct power hierarchy. Then, clearly $\Pi_{\beta}(\mathcal{A})\subset\mathcal{F}_{\beta}(\mathcal{A})$, which means there is a fragment of $\mathcal{F}_{\beta}(\mathcal{A})$ which is a proper extension of $\mathcal{A}$ that preserves any direct power sentence. Suppose, however, we would like to shape the cumulative power hierarchy itself, with all of its cross-level element functions, rather than a fragment of it, into a structure which preserves more first-order properties from its generating structure. That may be done in varying levels of faithfulness, in terms of satisfaction of $\mathcal{L}^{\sigma}$-sentences. The most faithful option, for successor levels, is to quotient each level $\mathcal{F}_{\beta^{+}}(\mathcal{A})$ by an adequate ultrafilter over $I_{\beta}$. But there is an intermediate level by quotienting $\mathcal{F}_{\beta}(\mathcal{A})$ with an appropriate equivalence relation, the product of which is isomorphic to $\Pi_{\beta}(\mathcal{A})$.
@@
Notice, furthermore, that following our considerations in Section 2, the above definition amounts to a hereditary definition of equality, that is, equality is defined modulo the $\vartheta$ function. It should therefore be no surprise the resulting structures are isomorphic to direct powers.
@@
(i) $\forall j\in I_{\rho(a)-1}\big(a(j)\equiv_{\rho(a)-1}b\big)$ and (ii) $\forall j\in I_{\rho(b)-1}\big(b(j)\equiv_{\rho(a)-1}c\big)$.
@@
either (iv) $a(k)=b$, or (v) $\forall j\in I_{0}\Big(\vartheta_{j}^{\delta}\big(a(k)\big)\equiv_{1}\vartheta^{\delta}_{t}(b)\Big)$.
@@
By (ii), we then get $a(k)\equiv_{1}c$. We conclude $a\equiv_{1}c$. The induction step is straightforward, as the analysis of each case goes to the previous level, where the induction hypothesis may be easily applied. ∎
@@
Let a direct power hierarchy and a cumulative power hierarchy of $\mathcal{A}$ both be generated and the family of index sets $\{I_{\alpha}\}_{\mathbf{On}}$. Then, for any $\beta\in\mathbf{On}$, $\Pi_{\beta}(\mathcal{A})\cong\mathfrak{F}_{\beta}(\mathcal{A})$.
@@
Therefore, cumulative powers may be seen as generalisations of direct powers, as the latter is obtained from the former by simply quotienting it by an equivalence relation defined by hereditary identity.
@@ 4. Ultrapowers and the hierarchy
As we have previously mentioned, to shape the hierarchy into the same first-order structure of its generating structure, it suffices to adapt the ultrapower construction to the present case. That would entail quotienting each level $\mathcal{F}_{\beta^{+}}(\mathcal{A})$ by an ultrafilter over $I_{\beta}$. The problem, of course, is that there are many elements in $\mathcal{F}_{\beta^{+}}(A)$, for $\beta>0$, which are not functions. More so, for $\beta>1$, there are pairs of elements who are functions, but whose domains differ. So simply quotienting $\mathcal{F}_{\beta^{+}}(\mathcal{A})$ by an ultrafilter over $I_{\beta}$ is not a valid construction. That could be worked around by quotienting $\mathcal{F}^{*}_{\beta^{+}}(A)$ by an ultrafilter over $I_{\beta}$. We follow a different route by defining the equivalence classes modulo an ultrafilter by using equality modulo the $\vartheta$ function, as in the construction of $\mathfrak{F}_{\beta}(\mathcal{A})$.
@@
Then, it is straightforward to adapt Łos’s theorem to the present case, so that we know:
@@
Therefore, quotienting any stage of the hierarchy by the $\sim$ relation produces a structure elementary equivalent to the one generating the hierarchy.
@@
Let $\mathcal{U}_{I}$ and $\mathcal{U}_{J}$ be ultrafilters over $I$ and $J$. The Rudin-Keisler ordering of ultrafilters $\leq_{\mathrm{RK}}$ is defined by $\mathcal{U}_{I}\leq_{\mathrm{RK}}\mathcal{U}_{J}$ when there is $h:J\to I$ such that $X\in\mathcal{U}_{I}$ iff $h^{-1}[X]\in\mathcal{U}_{J}$. Two ultrafilters are Rudin-Keisler equivalent, written $\mathcal{U}_{I}\approx_{\mathrm{RK}}\mathcal{U}_{J}$, when $\mathcal{U}_{I}\leq_{\mathrm{RK}}\mathcal{U}_{J}$ and $\mathcal{U}_{J}\leq_{\mathrm{RK}}\mathcal{U}_{I}$. [^12] In fact, that is only a preorder. The Rudin-Keisler (partial) ordering is defined on types of ultrafilters, which are defined as the equivalence classes of ultrafilters under the defined equivalence. For briefness, we use the concept as defined.
@@
Notice, if it was the case that any embedding between ultrapowers is of the form defined in Theorem 4.20, then given $|A|\leq|B|$, $\mathcal{U}_{I}\leq_{\mathrm{RK}}\mathcal{U}_{J}$ becomes a necessary and sufficient criterion for the embeddability of $\mathcal{A}^{I}/\mathcal{U}_{I}$ into $\mathcal{B}^{J}/\mathcal{U}_{J}$. Thus, we might raise our first open problem:
  • Open Problem 1: Given that $|A|<|B|$, is any embedding $e:\mathcal{A}^{I}/\mathcal{U}_{I}\to\mathcal{B}^{J}/\mathcal{U}_{J}$ of the form $f]_{\mathcal{U}_{I}}\mapsto[w(f)]_{\mathcal{U}_{J}}$ for some injective $u:\mathcal{A}\to\mathcal{B}$ and mapping $h:J\to I$ be a mapping? If not, is there a necessary and sufficient criterion of embeddability between $\mathcal{A}^{I}/\mathcal{U}_{I}$ and $\mathcal{B}^{J}/\mathcal{U}_{J}$?
@@
  • $\mathrm{R}_{j}^{\Upsilon_{\beta^{+}}(\mathcal{A})}[a_{1}]...[a_{n}]$ iff $\{x\in\Upsilon_{\alpha}(\mathcal{A})\mid\mathrm{R}_{j}^{\Upsilon_{\beta}(\mathcal{A})}a_{1}(x)...a_{n}(x)\}\in\mathcal{U}_{\beta}$, and for a limit $\lambda$, $\mathrm{R}_{j}^{\Upsilon_{\lambda}(\mathcal{A})}[a_{1}]...[a_{n}]$ iff $\mathrm{R}_{j}^{\Upsilon_{\delta}(\mathcal{A})}e^{\rho(a_{1})}_{\delta}(a_{1})...e^{\rho(a_{n})}_{\delta}(a_{1})$;
@@
Therefore, just as it was the case with direct powers, an ultrapower may be seen to be a structure emerging from quotienting a cumulative power by an appropriately defined equivalence relation.
@@
The above definitions generalize the notions of support, and its counterpart, of a structure, and of a function. $\mathrm{rep}(\mathcal{A})$ is divided following the reasoning that the elements of $\mathcal{A}^{I}/\mathcal{U}$ who are in its support should find representatives among the functions in the support of $\mathcal{A}^{I}$ (which are the functions whose support are their whole domains), and elements not in the support of $\mathcal{A}^{I}/\mathcal{U}$ should find representatives among the constant functions whose images are singletons of elements who are not in the support of their domain – intuitively, reflecting also the fact that each element not in the support of $\mathcal{A}^{I}/\mathcal{U}$ should be equivalent, in a sense, to some element in the anti-support of $\mathcal{A}$. In essence, the definition of $\mathrm{rep}(\mathcal{A})$ captures the intuition the ultrapower should not multiply the elements in the anti-support of its generating structure.
@@
(and likewise for $|S^{a\uparrow}_{\mathrm{F}}|$). Let $\mathrm{m}\in A$ the element such that $\tau_{\mathcal{A}}=\tau_{\mathrm{m}}=\sum_{\mathrm{F}\in\sigma}|S_{\mathrm{F}}^{\mathrm{m}\downarrow}|+|S_{\mathrm{F}}^{\mathrm{m}\uparrow}|$. We may notice that
@@
so since every ultrafilter is $\aleph_{0}$-complete, by we have our result. ∎
@@
The two above theorems offer the reason for which, in general, ultrapowers by non-principal ultrafilters of ordered structures, or structures with concurrent relations, do not embed into their related direct powers – for example, whereas Corollary 4.34 tells us a non-trivial ultrapower of the unordered naturals under the signature with solely the successor a function as a primitive symbol may embed into its relative direct power, Theorem 4.40 says the same structure with the addition of the usual total order, $\mathbb{N}_{\mathrm{PA},<}=\langle\mathbb{N}_{\mathrm{PA}},<\rangle$, is such that $\mathbb{N}_{\mathrm{PA},<}^{I}/\mathcal{U}\not\hookrightarrow\mathbb{N}_{\mathrm{PA},<}^{I}$ when $\mathcal{U}$ is free.
@@
    Let $e_{0,0}=\{h_{r_{0}},h_{r^{-1}_{0}}\}\cup\{h_{m}\mid m\in B\}$, where $h_{m}=\overline{m}$, $h_{r_{0}}:A\to L_{r_{0}}$ is some order preserving mapping such that $h_{r_{0}}(A)$ is cofinal in $L_{r_{0}}$, and $h_{r^{-1}_{0}}:A\to A$ is defined by:

### 18 Mar 2026 v2

@@ 2.2. Preservation and properties of cumulative powers
Therefore, for $n<\omega$, $\mathcal{F}^{+}_{n}(\mathcal{A})$ inherits any first-order property from $\mathcal{A}$ characterized by non-collapsible direct power sentences; and for $\beta\in\mathbf{On}$, $\mathcal{F}_{\beta}(\mathcal{A})$ inherits any first-order property from $\mathcal{A}$ characterized by constant-free non-collapsible direct power sentences – since, for $\beta>0$, $\mathcal{F}_{\beta}(\mathcal{A})$ does not have an interpretation of the constants of $\sigma$. Not only that, those are precisely the classes of first-order sentences inherited by cumulative powers. In other words, just as the fragment preserved by reduced products is precisely the one given by Horn sentences, and the fragment preserved by direct products is precisely the one characterized by Weinstein, the fragment preserved by (finite) cumulative powers is precisely the one given by non-collapsible direct power sentences.
Notice the restrictions of the results on non-collapsible formulas, in comparison to the formulas preserved by direct powers, concern precisely equality. That is related to the definition of the extended functions. In our former example, for $j\in\mathbb{N}$, $\vartheta^{1}_{j}(\overline{3})=\overline{3}(j)=3$, $\vartheta^{1}_{j}(\overline{-1})=-1$ and $\vartheta^{1}_{j}(2)=2$, so $\forall j\in\mathbb{N}\big(\vartheta^{1}_{j}(\overline{3})+\vartheta^{1}_{j}(\overline{-1})=\vartheta^{1}_{j}(2)\big)$. However, $\mathcal{F}_{1}(\mathbb{N})\not\models\overline{3}+\overline{-1}=2$, as $\overline{3}+\overline{-1}=\overline{2}\neq 2$. The restriction has also a deeper meaning. It concerns the difference between the definition of other extended relations and the equality relation. While at each level $\beta$ every other relation is defined by using the $\vartheta$ function, and therefore take into account all of the levels lesser than $\beta$ – so that, for example, if $\rho(a)<\rho(b)=\beta$ for a successor $\beta$, then $\mathrm{R}_{\beta}ab$ iff for each $j\in I_{\delta-1}$, $\mathrm{R}_{\beta-1}ab(j)$ –, the equality relation is interpreted as real identity, and thus holds between two objects iff they are the same object, regardless of how they are hereditarily related. In a few words, whereas every other relation is hereditarily defined, equality is not.
@@
Let $\mathcal{A}$ and $\mathcal{B}$ be $\sigma$-structures, and their cumulative power hierarchies be respectively generated by the families of index sets $\{I_{\alpha}\}_{\alpha\in\mathbf{On}}$ and $\{J_{\alpha}\}_{\alpha\in\mathbf{On}}$. If $\mathcal{A}\hookrightarrow\mathcal{B}$ and for each $\alpha\in\mathbf{On}$ $|I_{\alpha}|\leq|J_{\alpha}|$, then for any $\beta,\gamma\in\mathbf{On}$ such that $\beta\leq\gamma$, $\mathcal{F}_{\beta}(\mathcal{A})\hookrightarrow\mathcal{F}_{\gamma}(\mathcal{B})$.
@@ 3. Direct powers and the hierarchy
As the past results show, the cumulative power hierarchies generated, as a whole, present a structure that resembles those of their generating structures, but nevertheless fail to instantiate the generating structure itself. How can we recover more of the first-order properties of the generating structures? A way of doing that is by restricting the elements of the hierarchy to those of the parallel hierarchy which is not cumulative – in other words, a hierarchy of direct powers. Consider the following definition.
@@
Let $\{I_{\alpha}\}_{\alpha\in\mathbf{On}}$ be a family of index sets. The direct power hierarchy generated by $\mathcal{A}$ and $\{I_{\alpha}\}_{\alpha\in\mathbf{On}}$ is composed of the structures $\Pi_{\beta}(\mathcal{A})$, where each $\Pi_{\beta^{+}}(\mathcal{A})$ is the direct power of $\Pi_{\beta}(\mathcal{A})$ with index set $I_{\beta}$, and for a limit $\lambda$, $\Pi_{\lambda}(\mathcal{A})$ is the limit of embeddings of the preceding structures. More precisely:
@@
By Corollary 2.15, we know what sentences are preserved by $\Pi_{\beta}(\mathcal{A})$. Suppose we let the same family $\{I_{\alpha}\}_{\alpha\in\mathbf{On}}$ define both a cumulative power and a direct power hierarchy. Then, clearly $\Pi_{\beta}(\mathcal{A})\subset\mathcal{F}_{\beta}(\mathcal{A})$, which means there is a fragment of $\mathcal{F}_{\beta}(\mathcal{A})$ which is a proper extension of $\mathcal{A}$ that preserves any direct power sentence. Suppose, however, we would like to shape the cumulative power hierarchy itself, with all of its cross-level element functions, rather than a fragment of it, into a structure which preserves more first-order properties from its generating structure. That may be done in varying levels of faithfulness, in terms of satisfaction of $\mathcal{L}^{\sigma}$-sentences. The most faithful option, for successor levels, is to quotient each level $\mathcal{F}_{\beta^{+}}(\mathcal{A})$ by an adequate ultrafilter over $I_{\beta}$. But there is an intermediate level by taking a quotient of $\mathcal{F}_{\beta}(\mathcal{A})$ with an appropriate equivalence relation, the product of which is isomorphic to $\Pi_{\beta}(\mathcal{A})$.
@@
Notice, furthermore, that following our considerations in Section 2, the above definition amounts to a hereditary definition of equality, that is, equality is defined modulo the $\vartheta$ function. It should therefore be of no surprise that the resulting structures are isomorphic to direct powers.
@@
(i) $\forall j\in I_{\rho(a)-1}\big(a(j)\equiv_{\rho(a)-1}b\big)$ and (ii) $\forall j\in I_{\rho(b)-1}\big(b(j)\equiv_{\rho(b)-1}c\big)$.
@@
either (iv) $a(k)=b$, or (v) $\forall j\in I_{0}\Big(\vartheta_{j}^{\delta}\big(a(k)\big)\equiv_{1}\vartheta^{\delta}_{j}(b)\Big)$.
@@
By (ii), we then get $a(k)\equiv_{1}c$. We conclude $a\equiv_{2}c$. The induction step is straightforward, as the analysis of each case goes to the previous level, where the induction hypothesis may be easily applied. ∎
@@
Let a direct power hierarchy and a cumulative power hierarchy of $\mathcal{A}$ both be generated by the family of index sets $\{I_{\alpha}\}_{\mathbf{On}}$. Then, for any $\beta\in\mathbf{On}$, $\Pi_{\beta}(\mathcal{A})\cong\mathfrak{F}_{\beta}(\mathcal{A})$.
@@
Therefore, cumulative powers may be seen as generalisations of direct powers, as the latter is obtained from the former by simply taking a quotient of it by an equivalence relation defined by hereditary identity.
@@ 4. Ultrapowers and the hierarchy
As we have previously mentioned, to shape the hierarchy into the same first-order structure of its generating structure, it suffices to adapt the ultrapower construction to the present case. That would entail taking a quotient of each level $\mathcal{F}_{\beta^{+}}(\mathcal{A})$ by an ultrafilter over $I_{\beta}$. The problem, of course, is that there are many elements in $\mathcal{F}_{\beta^{+}}(A)$, for $\beta>0$, which are not functions. More so, for $\beta>1$, there are pairs of elements that are functions, but whose domains differ. So simply taking a quotient of $\mathcal{F}_{\beta^{+}}(\mathcal{A})$ by an ultrafilter over $I_{\beta}$ is not a valid construction. That could be worked around by taking a quotient of $\mathcal{F}^{*}_{\beta^{+}}(A)$ by an ultrafilter over $I_{\beta}$. We follow a different route by defining the equivalence classes modulo an ultrafilter by using equality modulo the $\vartheta$ function, as in the construction of $\mathfrak{F}_{\beta}(\mathcal{A})$.
The aim of the present section is to understand what issues which may arise from our considerations on the properties of cumulative powers may be inherited by them from ultrapowers. As such, more than showing how ultrapowers may be obtained from cumulative powers, it investigates the relation between a structure, its direct powers, and its ultrapowers by showing under what conditions: embeddability is lifted from given structures to their ultrapowers; each equivalence class of an ultrapower may find a representative among an intuitively defined subset of the unquotiented direct power; each equivalence class of an ultrapower may find a representative among an intuitively defined subset of the unquotiented direct power; an ultrapower may embed into a respective direct power of the same structure; an ordered field may embed into an ultrapower of a given subfield of it.
4.1. Obtaining elementarily equivalence from cumulative powers and embeddability between ultrapowers
@@
Then, it is straightforward to adapt Łoś’s theorem to the present case, so that we know:
@@
Therefore, taking a quotient of any stage of the hierarchy by the $\sim$ relation produces a structure elementarily equivalent to the one generating the hierarchy.
@@
Let $\mathcal{U}_{I}$ and $\mathcal{U}_{J}$ be ultrafilters over $I$ and $J$. The Rudin-Keisler ordering of ultrafilters $\leq_{\mathrm{RK}}$ is defined by $\mathcal{U}_{I}\leq_{\mathrm{RK}}\mathcal{U}_{J}$ when there is $h:J\to I$ such that $X\in\mathcal{U}_{I}$ iff $h^{-1}[X]\in\mathcal{U}_{J}$. Two ultrafilters are Rudin-Keisler equivalent, written $\mathcal{U}_{I}\approx_{\mathrm{RK}}\mathcal{U}_{J}$, when $\mathcal{U}_{I}\leq_{\mathrm{RK}}\mathcal{U}_{J}$ and $\mathcal{U}_{J}\leq_{\mathrm{RK}}\mathcal{U}_{I}$. [^13] In fact, that is only a preorder. The Rudin-Keisler (partial) ordering is defined on types of ultrafilters, which are defined as the equivalence classes of ultrafilters under the defined equivalence. For brevity, we use the concept as defined.
@@
Notice, if it were the case that any embedding between ultrapowers is of the form defined in Theorem 4.20, then given $|A|\leq|B|$, $\mathcal{U}_{I}\leq_{\mathrm{RK}}\mathcal{U}_{J}$ becomes a necessary and sufficient criterion for the embeddability of $\mathcal{A}^{I}/\mathcal{U}_{I}$ into $\mathcal{B}^{J}/\mathcal{U}_{J}$. Thus, we might raise our first open problem:
  • Open Problem 1: Given that $|A|<|B|$, is any embedding $e:\mathcal{A}^{I}/\mathcal{U}_{I}\to\mathcal{B}^{J}/\mathcal{U}_{J}$ of the form $f]_{\mathcal{U}_{I}}\mapsto[w(f)]_{\mathcal{U}_{J}}$ for some injective $u:\mathcal{A}\to\mathcal{B}$ and mapping $h:J\to I$? If not, is there a necessary and sufficient criterion of embeddability between $\mathcal{A}^{I}/\mathcal{U}_{I}$ and $\mathcal{B}^{J}/\mathcal{U}_{J}$?
@@ 4.2. Obtaining ultrapowers from cumulative powers
  • $\mathrm{R}_{j}^{\Upsilon_{\beta^{+}}(\mathcal{A})}[a_{1}]...[a_{n}]$ iff $\{x\in\Upsilon_{\beta}(\mathcal{A})\mid\mathrm{R}_{j}^{\Upsilon_{\beta}(\mathcal{A})}a_{1}(x)...a_{n}(x)\}\in\mathcal{U}_{\beta}$, and for a limit $\lambda$, $\mathrm{R}_{j}^{\Upsilon_{\lambda}(\mathcal{A})}[a_{1}]...[a_{n}]$ iff $\mathrm{R}_{j}^{\Upsilon_{\delta}(\mathcal{A})}e^{\rho(a_{1})}_{\delta}(a_{1})...e^{\rho(a_{n})}_{\delta}(a_{n})$;
@@
Therefore, just as it was the case with direct powers, an ultrapower may be seen to be a structure emerging from taking a quotient of a cumulative power by an appropriately defined equivalence relation.
4.3. Embedding an ultrapower into a direct power
@@
The above definitions generalize the notions of support, and its counterpart, of a structure, and of a function. $\mathrm{rep}(\mathcal{A})$ is divided following the reasoning that the elements of $\mathcal{A}^{I}/\mathcal{U}$ that are in its support should find representatives among the functions in the support of $\mathcal{A}^{I}$ (which are the functions whose supports are their whole domains), and elements not in the support of $\mathcal{A}^{I}/\mathcal{U}$ should find representatives among the constant functions whose images are singletons of elements that are not in the support of their domain – intuitively, reflecting also the fact that each element not in the support of $\mathcal{A}^{I}/\mathcal{U}$ should be equivalent, in a sense, to some element in the anti-support of $\mathcal{A}$. In essence, the definition of $\mathrm{rep}(\mathcal{A})$ captures the intuition that the ultrapower should not multiply the elements in the anti-support of its generating structure when they are definable.
@@
(and likewise for $|S^{a\uparrow}_{\mathrm{F}}|$). Let $\mathrm{m}\in A$ be the element such that $\tau_{\mathcal{A}}=\tau_{\mathrm{m}}=\sum_{\mathrm{F}\in\sigma}|S_{\mathrm{F}}^{\mathrm{m}\downarrow}|+|S_{\mathrm{F}}^{\mathrm{m}\uparrow}|$. We may notice that
@@
so since every ultrafilter is $\aleph_{0}$-complete, by Theorem 4.31 we have our result. ∎
@@
The two above theorems offer the reason for which, in general, ultrapowers by non-principal ultrafilters of ordered structures, or structures with concurrent relations, do not embed into their related direct powers – for example, whereas Corollary 4.34 tells us a non-trivial ultrapower of the unordered naturals under the signature with the successor function as its sole primitive symbol may embed into its relative direct power, Theorem 4.40 says the same structure with the addition of the usual total order, $\mathbb{N}_{\mathrm{PA},<}=\langle\mathbb{N}_{\mathrm{PA}},<\rangle$, is such that $\mathbb{N}_{\mathrm{PA},<}^{I}/\mathcal{U}\not\hookrightarrow\mathbb{N}_{\mathrm{PA},<}^{I}$ when $\mathcal{U}$ is free.
As the last point of this section, let $\mathbb{Q}^{*}$ be some construction of the hyperrationals, and $\infty$ and $\varepsilon$ be its infinite and infinitesimal elements, respectively. It is well known that the quotient ring $(\mathbb{Q}^{*}\setminus\infty)/\varepsilon$ is isomorphic to $\mathbb{R}$. We show now how that more generally applies between ordered fields and their subfields.
@@
    Let $e_{0,0}=\{h_{r_{0}},h_{r^{-1}_{0}}\}\cup\{h_{m}\mid m\in A\}$, where $h_{m}=\overline{m}$, $h_{r_{0}}:A\to L_{r_{0}}$ is some order preserving mapping such that $h_{r_{0}}(A)$ is cofinal in $L_{r_{0}}$, and $h_{r^{-1}_{0}}:A\to A$ is defined by: